log (225x2) = 3.456
Answer by stanbon(75887)

log (225x2) = 3.456
---
log(225x^2) = 3.456
225x^2 = 10^3.456
225x^2 = 2857.5905
x^2 = 12.7004
x = 3.5638 or x = -3.5648
=============================
Cheers,
Stan H.